OpenStudy (anonymous):

Solve K =1/3(b – 2a) for b

OpenStudy (across):

b=1/(3K)+2a

OpenStudy (anonymous):

how yu gt that

OpenStudy (across):

\[K=(1/(3(b-2a)))\]\[b-2a=1/3K\]\[b=1/3K+2a\]

OpenStudy (anonymous):

it is (1+2ak)/3k

OpenStudy (across):

Although I can't tell if 1/3 is multiplying (b-2a) or not.

OpenStudy (anonymous):

xtremely sorry ......ans is (1+6ak)/3k